A rocket is launched straight up from the ground with constant acceleration. When it reaches a height of 240m, a bolt falls off the side of the rocket. The bolt hits the ground 6.0s later.

A) How fast was the bolt (or rocket) going when it broke loose from the rocket?

B) What was the rocket's acceleration?

C) What was the maximum height that the bolt reached?

Thanks.
Answer #1

a)

using 2nd equation of motion

h = - ut + 0.5 gt^2

- 240 = - 6u + 4.9* 6^2

u = 69.4 m/s

======

b)

u^2 = 0 + 2* a * 240

a = 10.03 m/s^2

======

c)

max height reached by bolt

h' = h + u^2/(2g)

h' = 240 + 69.4^2 / (2 *9.8)

h' = 485.733 m

=====

do comment in case any doubt, will reply for sure . Goodluck